# Behavioral Manifolds: Representing the Landscape of Grasp Affordances in the Relative Pose Space

### v1

@@ I Introduction
Empirical methods typically emulate previously established grasp strategies. This is commonly pursued by recording human grasps, and applying RL techniques to emulate them with the considered robotic system. The range of possible grasps are defined in an established taxonomy ([13]). In [14], the authors use a virtual reality setup to train a robot to imitate human behavior during task execution. Supervised learning has also been used to support grasp synthesis, either by observing and reproducing successful grasps ([15]), or via an annotated mix of failed and successful grasps ([16]). Recent approaches use meta-learning to pursue one-shot imitation capabilities ([17, 18]) from visual input. These models are trained using general grasp training sets (such as those in [19, 20]), then fine-tuned via demonstration. Other empirical approaches have developed grasp synthesis strategies by applying reinforcement learning techniques to physical setups featuring robotic arm, end-effector, and a series of objects of interest. The approach amounts to repeatedly attempting to grasp an object, and iteratively improving grasping behavior. Other empirical approaches rely on template matching, whereby a set of template grasps are defined for a considered set of objects. When a novel object is encountered, its shape is analyzed, similarities to known shapes identified, and an appropriate template grasp is selected. This is typically pursued by either mapping template reference frames of grasps to new objects ([21, 22]), or by simulating a considered grasp before execution ([21]).
@@
To address these issues, we propose a generative approach to learning grasp affordances. We rely on the existence of a grasp synthesis policy which, given an admissible relative pose between gripper and object, defines what trajectories are applied to the gripper’s degrees of freedom (gripper base and actuated axes or joints) in pursuit of grasp emergence. Such a policy $\mathcal{G}$ provides a map from relative pose space (denoted $\mathcal{P}$ in the following) to grasp configuration space $\mathcal{C}$; that is, $\mathcal{G}:\mathcal{P}\rightarrow\mathcal{C}$. Relying on the existence of such a map, one is able to approach the problem of learning how to grasp an object by formulating grasp policies over the relative pose space $\mathcal{P}$, rather than the configuration space $\mathcal{C}$, which affords a number of benefits. Under consistent grasp policy, well-defined sets in $\mathcal{P}$ can be found to lead to consistent grasp types or behaviors. Establishing the landscape of such behavioral manifolds (i.e. sets of points $p\in\mathcal{P}$ mapping to consistent grasp types in $\mathcal{C}$ under $\mathcal{G}$, see example in Fig. 1) provides an explicit representation of the range of grasp solutions for the gripper, object, grasp policy tuple.
In the following, we explore the merit of the proposed approach. In Section II, we discuss the use of a previously introduced dynamic quality metric ([26]) to assess grasp resilience during object manipulation. Then, in Section II, we propose a simple example of grasp synthesis policy, and use it to establish the behavioral landscape in relative pose space, for a number of example objects. We further assess impact on performance of informing grasp synthesis using haptic information, and evaluate impact in terms of the size of behavioral manifolds. Section IV concludes this paper.
@@ III-D Grasp Policies Comparison
IV Conclusion
@@
In this paper, we proposed a novel approach to grasp affordance learning, approaching the problem from the relative pose space. The method allows visualisation of possible grasp outcomes (both grasp types and quality) for different grasp synthesis policies. Different approaches to grasp synthesis are discussed, emphasizing the use of shape adaptation, supported by a haptic modality. Current efforts are investigating efficient numerical implementation of the approach. Preliminary results show convergence times 4 to 6 times faster (depending on grasped object complexity) than those observed using PPO on identical hardware. In addition, the approach is being extended to address assembly and disassembly problems, accounting from transmission of efforts, through the grasped object, onto an assembly.

### v2

# Systematically Exploring the Landscape of Grasp Affordances via Behavioral Manifolds

@@ I Introduction
Empirical methods typically emulate previously established grasp strategies. This is commonly pursued by recording human grasps, and applying RL techniques to emulate them with the considered robotic system. The range of possible grasps are defined in an established taxonomy ([13]). In [14], the authors use a virtual reality setup to train a robot to imitate human behavior during task execution. Supervised learning has also been used to support grasp synthesis, either by observing and reproducing successful grasps ([15]), or via an annotated mix of failed and successful grasps ([16]). Recent approaches use meta-learning to pursue one-shot imitation capabilities ([17, 18]) from visual input. These models are trained using general grasp training sets (such as those in [19, 20]), then fine-tuned via demonstration. Other empirical approaches have developed grasp synthesis strategies by applying reinforcement learning techniques to physical setups featuring robotic arm, end-effector, and a series of objects of interest. The approach amounts to repeatedly attempting to grasp an object, and iteratively improving grasping behavior. Other empirical approaches rely on template matching, whereby a set of template grasps are defined for a considered set of objects. When a novel object is encountered, its shape is analyzed, similarities to known shapes identified, and an appropriate template grasp is selected. This is typically pursued by either mapping template reference frames of grasps to new objects ([21, 22]), or by simulating a considered grasp before execution ([21]). Recently, new developments have focused on artificial expansion of existing image datasets for zero-shot learning. In [23], the authors use a large language model, coupled with a latent diffusion model, to generate a large dataset of objects for grasping. The images are automatically annotated to identify viable grasps using a prismatic gripper. While this approach does create a large dataset, the annotations lack real-world information about the included objects; the generated images lack depth, and vital parameters of created objects such as mass, inertia, and material friction coefficients are missing.
@@
Other approaches focus on more complex object geometries. In [27, 28], the authors propose a method of contact-invariant optimization that enables the generation of grasp trajectories designed to achieve a specific task. Later on, Grasp’d adapts this approach to generate grasps for a given gripper-object pair ([29]). An alternative approach to grasp generation is presented by the authors in [30]. They propose Multi-fingan, an algorithm to generate grasps from RGB-D information. By training a classification network on manually generated grasps, they are able to rapidly infer new grasps for presented objects. However, these approaches aim to find and classify a single grasp for a given gripper-object pair. In case of a changed scenario, the grasp must be regenerated.
To address the aforementioned issues, we propose a generative approach to learning grasp affordances. We rely on the existence of a grasp synthesis policy which, given an admissible relative pose between gripper and object, defines what trajectories are applied to the gripper’s degrees of freedom (gripper base and actuated axes or joints) in pursuit of grasp emergence. Such a policy $\mathcal{G}$ provides a map from relative pose space (denoted $\mathcal{P}$ in the following) to grasp configuration space $\mathcal{C}$; that is, $\mathcal{G}:\mathcal{P}\rightarrow\mathcal{C}$. Relying on the existence of such a map, one is able to approach the problem of learning how to grasp an object by formulating grasp policies over the relative pose space $\mathcal{P}$, rather than the configuration space $\mathcal{C}$, which affords a number of benefits. Under consistent grasp policy, well-defined sets in $\mathcal{P}$ can be found to lead to consistent grasp types or behaviors. Establishing the landscape of such behavioral manifolds (i.e. sets of points $p\in\mathcal{P}$ mapping to consistent grasp types in $\mathcal{C}$ under $\mathcal{G}$, see example in Fig. 1) provides an explicit representation of the range of grasp solutions for the gripper, object, grasp policy tuple. Additionally, we propose a systematic method of exploring grasp configuration space via relative pose space. Using this approach, we can quickly generate a concise mapping that describes regions with similar grasp configurations. This enables us to quantify and compare the performance of reachable grasp configurations.
In the following, we explore the merit of the proposed approach. In Section II, we discuss the use of a previously introduced dynamic quality metric ([31]) to assess grasp resilience during object manipulation. Then, in Section II, we propose a simple example of grasp synthesis policy, and use it to establish the behavioral landscape in relative pose space, for a number of example objects. We further assess impact on performance of informing grasp synthesis using haptic information, and evaluate impact in terms of the size of behavioral manifolds. In Section IV, we showcase our proposed method for systematically exploring the relative pose space to efficiently find and categorize regions with similar grasp configurations. Section V compares our approach to existing end-to-end learning methods. Section VI concludes this paper.
@@ III-D Grasp Policies Comparison
IV Grasp Space Exploration
Figure 6: 2D-cutout of grasp synthesis regions for a cup, generated with a Robotiq 3-fingered gripper. The red line shows the path Google’s PPO-based space exploration algorithm takes, with green arrows denoting jumps of search paths and yellow dots indicating the individual maximas found in each region. The right graph shows a comparison of execution time between PPO and our adaptive search.
In section III, we describe how to synthesize grasps for a given object. Section II introduces a metric to evaluate a grasp’s performance.In this section, we will expand on the introduced methods to generate an encompassing set of grasps for an object that covers most approach trajectories, while also ranking said grasps according to their performance. This necessitates a method of exploring grasp space. As previously discussed in section III however, purely searching through all possible grasp configurations requires an unreasonable amount of time. Instead, we explore the lower-dimensional set of relative base poses $\mathcal{P}$ and their mapped grasp configuration set $\mathcal{C}$. This enables us to discard unreachable grasp configurations and focus on a reasonable subset of valid, usable grasps.
IV-A PPO-based Grasp Space Exploration
Our initial approach was based on using a Proximal Policy Optimization (PPO)-based reinforcement learning approach, as presented in [45]. PPO algorithms have previously been used for successful grasp policy learning, as shown by the approaches demonstrated in [46, 47, 48, 49]. We built our setup on top of the AngoraPy framework for goal-oriented research ([50]). The framework facilitates a method of integrating various physics simulators and connecting them to reinforcement learning algorithms. For our use-case, we developed a custom physics simulator based on Bullet that allows us to control a gripper, evaluate its interaction with various objects, and compute a grasp affordance metric as presented in Section II. AngoraPy then extends an API to connect this simulator with a PPO algorithm to learn grasp policies via reinforcement learning. However, as PPO is a model-free algorithm, we must still provide an our own input and output. Here, our goal is to explore as many grasps as possible for a given object. As stated previously, we are indirectly covering the grasp configuration space $\mathcal{C}$ by exploring the relative pose space $\mathcal{P}$. Our aim is to determine the most stable grasp while minimizing the employed grasp force. Taking both goals into account, our input for a single PPO cycle(?) is the initial relative gripper pose $\eta_{\rm rp}$ as well as the maximum allowed gripper force. The output of the simulation is our grasp metric $\mu_{\rm s}$. We configure the simulator’s initial conditions by adding both the considered object as well as our gripper at its relative pose. We then run the simulation and synthesize a grasp as described in section III. If the synthesis fails to generate a grasp, we mark the initial pose as invalid and return a score of 0. If the synthesis is successful, we record the resultant grasp configuration $(\eta_{\rm rc},\theta_{\rm g})$ in our set of valid grasp configurations $\mathcal{C}$. Lastly, we simulate the given trajectory, and return the computed grasp metric score $\sigma_{\rm s}$ (should the grasp fail during execution, the score is 0). Figure 7 shows an example of the individual steps.
We define the PPO loss function as
$L=w_{\rm s}\mu_{\rm s}+w_{\rm f}\tau_{\rm m},$ (12)
where $w_{\rm s}\in\mathds{R}^{+}$ is the weight assigned to the quality of the evaluated grasp, $\tau_{\rm m}\in\mathds{R}^{+}$ is the maximum joint torque employed by the evaluated grasp during trajectory execution in Nm, and $w_{\rm f}\in\mathds{R}^{+}$ is weight assigned to $\tau_{\rm m}$ in $\frac{1}{\rm Nm}$.
An example 2D-cutout of a PPO-based exploration is shown in Fig. 6. The search begins at the top left and randomly explores until it finds a region of valid poses. Once achieved, it performs a gradient descent until a local optimum is found, then continues it’s search. The approach covers most of the pose space around the mug and finds the pose regions described above. Later, we can group the poses into their respective regions (labeled $\mathcal{A}$$\mathcal{D}$) by looking at their mapped grasp configurations. With this, we are able to determine a good grasp for a given object within around 20-25 minutes. However, during the development of our approach, we noted a series of properties that could be used to improve search times. However, PPO is a model-agnostic optimization method. It does not require information about the underlying system. Therefore, while it can be adapted to a multitude of problems without too much effort, it also cannot take advantage of system characteristics while performing a search. To properly use our findings, we had to adapt a separate approach.
IV-B Adaptive Grasp Space Exploration
We want to perform a comprehensive exploration of reachable grasp configurations PPO-based approach searches for the best grasp using a heuristic approach. We aim to use the object’s geometric properties to perform a systematically search. By leveraging our behavioral mapping, we can investigate the simple pose space instead of the more complex and higher-dimensioned configuration space. By exhaustively searching through relative pose space, we can cover a reasonable set of grasp configurations. We propose a succinct search algorithm that covers nearby pose space and focuses on areas of interest (regions where a mapping to a valid grasp configuration exists). Identify and characterize regions within relative pose space that map to similar grasp configurations. Use spherical coordinates to define the pose of the gripper base’s CoM in relation to the object’s CoM. The most important aspect of a mapping is the approach vector. By using spherical coordinates, we can use $\theta,\phi$ to define approach vector and use radius $r$ to adjust distance to nearest object point. First iterate over both $\theta$ and $\phi$ in steps of $\Delta\theta$. At each position, adjust $r$ so that the initial distance between the gripper base’s surface is at a distance of $r_{\rm i}$ from the closest object point. After setting the gripper base’s relative position, we adjust its attitude so that the gripper base’s surface normal is aligned with the normal of the closest surface point on the object. We also iterate over the same $\theta$, $\phi$, but with a variation in $r$ from $r_{\rm min}$ to $r_{\rm max}$ by $\Delta r$. At each position, we check whether the pose can be mapped to a valid grasp configuration. Grasp configurations are identified, and poses that map to similar grasps are added to individual lists. Using this initial pass-through, we attain a rough outline of possible grasp regions. Next, we perform a search to determine the border of detected regions. We check our results for existing poses indicating a nearbyg border of a region (Coordinates where one pose can be mapped to a valid grasp, and a neighboring coordinate without this property). For each of these coordinates, we performa a search over closer neighboring points with a smaller step size of $\Delta\theta/2$ and $\Delta r/2$. Each new point that is identified to be within the region is added to the list. We continue iterating over region coordinates until the slopes of line segments between neighboring poses are smaller than $s_{\rm max}$. This enables us to determine regions with arbitrary accuracy. Within each region, we perform a gradient descent search to find the pose with best grasp quality.
The following lists our observations, then describes how our approach is configured. 1. Multiple poses in $\mathcal{P}$ map to the same or very similar grasps in $\mathcal{C}$. If an exploration algorithm is aware of this, it can infer the grasp metric score based on previous evaluation of similar grasps. This would allow us to skip a simulation step, saving time. 2. Initial poses along a local normal of an object surface map to the same grasp configuration. Usually, grasps don’t change when approaching from the same direction. An example of this is shown in Fig. 6, in set $\mathcal{A}$. Here, all poses along the object surface normals are mapped to the same grasp. By identifying regions where this is the case, we can skip both grasp synthesis and evaluation along these normals, as all poses will result in the same grasp. 3. Exploring objects with shape symmetries result in grasp regions with similar symmetries. For example, the cup shown in Fig. 6 is mirror symmetric along the horizontal axis, and the pose region mirror this symmetry. Taking these properties into account, we created our own search method, replacing the PPO-based approach; not only does our method speed up the required search time, it also returns more information. Instead of only searching for the most performant grasp, we generate a map of all possible approach trajectories and their corresponding grasp configuration, along with a ranking of each grasp.
Exploring such a large space in a reasonable amount of time is difficult. Therefore, we developed a method to simplify configuration space exploration; we limit the search to all possible relative base pose, and generate one reasonable joint configuration per pose. Similar approaches have been used in . The resulting exploration space is reduced to a 6-dimensional manifold (?) consisting of all possible relative base poses, removing the need to explore multiple joint configurations for each individual pose. The PPO performs a search through this configuration space to find the base pose resulting in the best grasp metric. In the following, we describe both our approach to generating a reasonable grasp joint configuration for a given relative gripper base pose as well as how we compute the grasp affordance metric for this grasp. See Fig. 7 for an example of the individual phases of a single grasp evaluation cycle.
1. Initialization
3. Trajectory Execution and Grasp Analysis
2. Grasp Synthesis
Figure 7: Evaluation of a relative pose. Top left image displays the initial gripper configuration, with the gripper base placed at an initial pose relative to the object, top right image shows the synthesized grasp, and bottom image shows the execution of an example trajectory.
IV-C Performance Comparison
Comparing both approaches, we note a significant reduction in search times. Examples are shown in the graph of Fig. 6. For both simple as well as complex shapes, we note a significant reduction in exploration time. Additionally, the adaptive approach yields a complete mapping of nearby pose to grasp configuration space, with each grasp tagged with a corresponding quality metric. PPO-based exploration can yield similar results, but this is not guaranteed; it’s main objective is finding the most performant grasp while still sufficiently exploring all possible options.
We evaluated our approach on the YCB dataset and compared it to using a conventional PPO-based approach. We evaluated our approach while testing various number of attitudes per position (1,2,4,8 rotated around the roll axis of the gripper’s base, and 1,2,4 around the pitch and jaw axis). When using our approach with only a single orientation per test position, we note a significant decrease in required search times. However, we can also show that the number of evaluated points is far lower, and we miss many possible grasp configurations. Increasing the number of tested attitudes when rotating around the roll axis up to 4 also increases exploration time, but greatly increases the number of found valid poses. Increasing to 8 no longer significantly valid poses, but still nearly doubles search times. Table 8 shows how the number of sampled attitudes per position compares to PPO performance over the entire YCB dataset. Conversely, increasing the number of tested rotations around the pitch and jaw axes does not significantly increase the number of found valid poses. This is due to the fact that most of these grasp configurations have already been discovered by testing other relative base positions. Due to the fact that each relative position is assigned an attitude pointing to the closest surface point, rotating around the pitch/jaw axes merely results in a grasp synthesis that targets an object surface that’s already covered by a different initial relative position.
Att. Quality # of Grasps Misses Time
1 76.42% 67.71% 94.48% 23.61%
2 83.91% 82.48% 84.68% 42.81%
4 99.31% 98.32% 76.46% 73.52%
8 100.03% 100.61% 75.16% 170.19%
Figure 8: Comparison of our approach with PPO. Rows shows the performance of our aproach when evaluating amounts of attitutes per position as a fraction of PPO performance, averaged over the YCB dataset. On average, PPO achieved a quality 0.78, with 6234.53 grasps identified, 4527.64 misses, over 34.65min per object.
V Comparison to End-to-End Learning Methods
As mentioned in Section I, many recent approaches to grasp generation focus on deep learning methods that combine the entire pipeline into a single algorithm. As many components are closely interconnected, this approach has shown to be effective in rapidly generating grasps for novel objects. In the following, we briefly compare our approach as outlined in the previous sections to two end-to-end methods (Multi-fingan [30] and Grasp’d [29]), showcasing advantages and disadvantages of both approaches. The first considered aspect is grasp generation time. Here, established approaches can generate a grasp in real-time, whereas our approach requires at least a minute to find a suitable grasp for a considered object. However, once a complete pose-grasp mapping is generated, the required time is drastically reduced down to approximately 10s (most likely, this could be further reduced by improving our search algorithm). Regardless, established methods are faster. The next aspect is adaptability. All approaches can handle arbitrary object shapes. However, our approach is held general and can easily be adjusted to other gripper types, regardless of whether it contains prismatic or revolute joints. End-to-end based approaches must be retrained for each new gripper type, which necessitates the generation of demonstrators or supervised learning setups, depending on the approach used. Either options requires a considerable investment of time and resources. Additionally, we can show that our approach generates better grasps. When compared over the YCB dataset, we consistently outperform comparable methods, as shown in table 3. As shown, this is consistent for various metrics, even those we were not actively monitoring during exploration. Lastly, we considered expended gripper effort. We compared which grasps all approaches generate and how much joint effort is required to maintain an established grasp while holding an object under normal gravity conditions. Table 4 shows the results; as can be seen, the required efforts remain comparable when considering simple shapes (e.g. balls, cubes, …). However, when generating grasps for shapes containing ridges/ledges, requires less effort to maintain its grasp. This is due to the fact that our method takes indirect efforts into account, e.g. ledges being in contact with digit surfaces. An example is shown in Figure TODO. For this, all algorithms generated grasps on a power drill from the same initial reference pose. Our algorithm created a grasp that has the drill’s motor section resting on the top surface of the gripper, whereas other approaches grasp without taking advantage of this shape feature. As a result, the end-to-end-generated grasps must compensate for gravity via exerting a sufficient amount of friction forces, whereas our grasp passively counters gravity. Similar results can be seen for most complex shapes from the YCB-dataset, leading to end-effectors requiring 4.6% less joint efforts when employing grasps generated using our method. Note that, as mentioned above, required efforts are negligibly different when considering simple shapes without ridges.
Figure 9: Robotiq end-effector grasping a screwdriver. Due to it’s shape, only one valid grasp type is available.
Quality Metric Our Approach Multi-fingan Grasp’d
$\mu_{\rm s}$ 0.86 0.74 0.75
$\mu_{\rm g}$ 0.96 0.94 0.94
$\mu_{\rm f}$ 0.78 0.73 0.72
$\mu_{\rm m}$ 0.83 0.79 0.78
Table 3: Average quality of grasps generated by end-to-end methods and our approach over the YCB dataset.
$\tau_{\rm s}$ $\tau_{\rm eg}$ $\tau_{\rm em}$
Simple Shapes
Sphere 1.48Nm 1.48Nm 1.48Nm
Cube 1.63Nm 1.64Nm 1.63Nm
Dumbbell 1.29Nm 1.53Nm 1.29Nm
Pringles Can 1.54Nm 1.54Nm 1.55Nm
Small Objects
Spoon 0.83Nm 0.84Nm 0.81Nm
Fork 0.83Nm 0.84Nm 0.82Nm
Screwdriver 0.93Nm 0.99Nm 1.00Nm
Uneven Mass
Power Drill 1.87Nm 2.13Nm 2.31Nm
Hammer 2.32Nm 2.32Nm 2.30Nm
Delicate Objects
Banana 0.52Nm 0.52Nm 0.53Nm
Tomato 0.51Nm 0.53Nm 0.53Nm
Foam Ball 0.20Nm 0.20Nm 0.20Nm
Textured Surface
Stanford Bunny 0.88Nm 0.95Nm 0.95Nm
Spiked Ball 1.46Nm 1.46Nm 1.46Nm
Table 4: Required joint efforts required to maintain a generated grasp. We compare the grasps generated using our approach ($\tau_{\rm s}$), Multi-fingan ($\tau_{\rm eg}$), and Grasp’d ($\tau_{\rm em}$).
VI Conclusion
@@
In this paper, we proposed a novel approach to grasp affordance learning, approaching the problem from the relative pose space. The method allows visualisation of possible grasp outcomes (both grasp types and quality) for different grasp synthesis policies. Different approaches to grasp synthesis are discussed, emphasizing the use of shape adaptation, supported by a haptic modality. Using a systematic approach to pose space exploration enables a convergence time Preliminary results show convergence times 33% faster than those observed using PPO on identical hardware. In addition, the approach is being extended to address assembly and disassembly problems, accounting from transmission of efforts, through the grasped object, onto an assembly.